# Finding Safety Violations of AI-Enabled Control Systems through the Lens of Synthesized Proxy Programs

### v1

@@ 7.4. Handling Large AI Controllers
As mentioned in Section 5.1, the AI controllers used in our evaluations vary significantly in size, with neuron counts ranging from 440 to 1200, which are relatively large compared to those used in other studies like (Zhang et al., 2022; Tran et al., 2020), providing a diverse basis for evaluation. However, we anticipate that larger models will be integrated into control systems as hardware continues to advance. Therefore, we conduct additional experiments to evaluate Synthify’s performance on control systems with larger AI controllers. These new experiments use three variant models trained in the CartPole, Pendulum, and Self-Driving control systems, also released by Zhu et al. (Zhu et al., 2019), which feature larger AI controllers with 3 layers containing 1200, 900, and 800 neurons, respectively—an increase of 2.4$\times$ to 6.6$\times$ the number of neurons compared to the 8 controllers used in our main experiments.
@@ 7.7. Potential Applications in End-to-End AI Systems
First and foremost, Synthify enhances the efficiency and sub-specification coverage of falsification for end-to-end AI systems. As mentioned in Section 1, these systems typically involve complex AI controllers with numerous parameters for processing inputs and generating control signals in one go (Bojarski et al., 2016; Julian et al., 2016). This complexity leads to high computational costs during testing (Haq et al., 2022; Pang et al., 2022). Synthify mitigates this by synthesizing efficient proxy programs that replace the AI controllers, significantly reducing the runtime of falsification. Moreover, thorough testing of end-to-end AI systems requires diverse test inputs (Lou et al., 2022; He et al., 2024). Synthify meets this need by maximizing sub-specification coverage of the conjunctive safety specification, as a more diverse set of test inputs is likely to falsify more sub-specifications (Zhang et al., 2023; Viswanadha et al., 2021); intuitively vice versa. The $\epsilon$-greedy sampling strategy employed by Synthify promotes the exploration of various sub-specifications, enabling the falsification process to uncover diverse safety violations by targeting different aspects of the safety requirements.
@@
Furthermore, certain modules in Synthify, such as our synthesis and $\epsilon$-greedy sampling strategy, can complement existing verification and testing tools for end-to-end AI systems. For instance, our synthesis module could be applied to generate simple programs that closely approximate the behavior of complex AI controllers but are more amenable to formal verification methods (Fulton et al., 2020; Landers and Doryab, 2023), like constructing barrier certificates (Zhu et al., 2019) or performing reachability analysis (Tran et al., 2019). In the context of testing, our $\epsilon$-greedy sampling strategy can be integrated into existing falsification tools like (Dreossi et al., 2019a; Zhang et al., 2022) to strengthen their ability to falsify conjunctive specifications. Alternatively, it can support automated testing frameworks that handle multiple objectives (Panichella et al., 2018; Haq et al., 2022), helping developers efficiently prioritize and focus on the most promising objectives during testing.
@@
Beyond testing and verification, Synthify aids developers in inspecting end-to-end AI systems for interpretability and explainability—two essential factors in ensuring the trustworthiness of AI-enabled systems (Glanois et al., 2024). Interpretability focuses on understanding the system’s inherent characteristics. Developers can leverage Synthify’s identified safety violations and corresponding falsifying inputs to derive formal specifications and properties of AI controllers (Gopinath et al., 2019; Rozier, 2016), serving as proof of the system’s safety (or lack thereof). Explainability, on the other hand, involves using external proxies to clarify the system’s operations (Vouros, 2022). Synthify’s synthesized proxy programs, with straightforward control logic, can provide explanations for generated control signals by transparently revealing the execution process. This allows developers to better understand the AI controller’s decision-making and identify potential vulnerabilities or biases within the system.
@@ 8.1. Falsification of AI-Enabled Control Systems
Falsification (Eddeland et al., 2020; Annpureddy et al., 2011; Adimoolam et al., 2017; Menghi et al., 2020; Waga, 2020; Ernst et al., 2021b; Tuncali et al., 2018; Chandratre et al., 2023) has emerged as an alternative to formal verification that provides better efficiency to uncover violations of a given safety specification. There comes many falsification techniques supported by optimization algorithms, such as Hill Climbing (Mathesen et al., 2021), Simulated Annealing (Abbas and Fainekos, 2012; Thibeault et al., 2021), Monte Carlo Tree Search (Zhang et al., 2018, 2021b), gradient-based methods (Abbas et al., 2014; Yaghoubi and Fainekos, 2017), as well as some machine learning-guided methods (Aineto et al., 2023; Zhang et al., 2021a).
To date, there has been limited attention directed toward the falsification of AI-enabled control systems, with only a few early attempts (Yaghoubi and Fainekos, 2019; Song et al., 2022; Zhang et al., 2022; Dreossi et al., 2019b; Xie et al., 2023). Among these studies, Song et al. (Song et al., 2022) evaluate the reliability of AI-enabled control systems using two existing falsification tools (Annpureddy et al., 2011; Donzé, 2010) and highlight that they need further improvements, which necessitates our work. Zhang et al. (Zhang et al., 2022) and Yaghoubi et al. (Yaghoubi and Fainekos, 2019) develop gray-box falsification techniques that make different assumptions compared to our study. They assume the AI controllers or plants accessible and collect not only the input and output signals from the control system but also intermediate information of the AI controllers or plants as guidance for falsification. In comparison, our approach is a black-box falsification tool that solely relies on the input and output signals, making it more practical particularly in real-world applications where the controller or plants can be complex or even inaccessible if they are confidential products from third-party vendors.
@@ 8.2. Surrogate-Assisted Falsification
Another set of studies integrates surrogate models into the falsification process. For instance, several studies (Deshmukh et al., 2017; Ramezani et al., 2022; Moss et al., 2023; Zhang and Arcaini, 2021; Qin et al., 2022) employ Bayesian optimization with Gaussian processes as surrogate models to estimate the robustness semantics of STL specifications, guiding the search for safety violations. Additionally, Peltomäki et al. (Peltomäki and Porres, 2022) propose training a surrogate DNN model to directly generate falsifying inputs for control systems.
While our synthesized proxy programs may resemble surrogate models, they fundamentally differ from those in above studies. First, in terms of form, our proxy programs implement linear controllers, whereas surrogate models in prior work are typically DNNs (e.g., (Kundu et al., 2024; Peltomäki and Porres, 2022)) or machine learning models (e.g., (Deshmukh et al., 2017; Qin et al., 2022)). The simplicity of our programs offers greater efficiency. Second, regarding purpose, our proxy programs replicate the functionality of AI controllers, acting as their stand-ins to facilitate efficient falsification by minimizing the need to run the controllers themselves. This addresses the unique computational challenges associated with AI controllers, which are expensive to execute due to their vast number of parameters. In contrast, surrogate models in earlier work are usually designed to replace control systems or predict robustness semantics. Third, in terms of underlying technique, our approach utilizes Evolution Strategy for program synthesis, while prior methods often rely on Bayesian optimization (e.g., (Qin et al., 2022; Ramezani et al., 2022)) or DNNs (e.g., (Menghi et al., 2020; Kundu et al., 2024; Peltomäki and Porres, 2022)) to build surrogate models.
@@ 8.3. Program Synthesis for AI models
Although program synthesis has been studied for decades (Sobania et al., 2022), very little attention has been paid to program synthesis specialized for AI models. A few studies (Verma et al., 2018; Trivedi et al., 2021) have proposed methods for generating programmatic policies from well-trained DRL models, which are more interpretable than neural networks. Additionally, Zhu et al. (Zhu et al., 2019) introduce an inductive synthesis-based method that can synthesize deterministic and verified programs from DRL-based models. Inala et al. (Inala et al., 2020) and Qiu et al. (Qiu and Zhu, 2021) synthesize programmatic controllers to perform continuous control tasks. Our approach differs from these studies in that we propose an Evolution Strategy-based algorithm that can effectively synthesize programs as proxies for AI controllers, which we believe is both technically novel and offers a unique perspective on improving the efficiency of falsification in AI-enabled control systems.
@@ 8.4. Multi-Objective Optimization in AI-Enabled Control Systems
In the context of testing, Zolfagharian et al. (Zolfagharian et al., 2023) and Haq et al. (Haq et al., 2022, 2023) propose multi-objective search-based testing methods for AI-enabled control systems, accounting for diverse fitness functions or safety requirements. Further studies address repairing these systems. Lyu et al. (Lyu et al., 2023) repair AI controllers’ unsafe behaviors by adding patches to control signals via multi-objective optimization, while MORTAR (Wang et al., 2024) replaces unsafe signals from AI controllers with repaired alternatives through gradient-based optimization. Both methods solve the multi-objective problem of maximizing satisfaction of safety requirements while minimizing changes to the original control signals. Regarding safety enhancement without directly repairing the violations, SIEGE (Song et al., 2023) stands out by presenting an ensemble framework that synergistically combines multiple AI controllers to improve the system’s reliability across multiple control objectives. GenSafe (Zhou et al., 2024b) corrects the behaviors of AI controllers that potentially deviate from safety constraints during training to enhance their safety while also maintaining performance at a satisfactory level. In addition to all of the above, Zhou et al. (Zhou et al., 2024a) establish a benchmark for multiple robotic manipulation tasks, providing a platform to evaluate the safety and effectiveness of AI controllers in a multi-objective context.
@@ 9. Conclusion and Future Work
Replication Package. For reproducibility, we have made our code and data publicly available: https://figshare.com/s/0c7fd68fd62cd5ecfd8c.

### v2

@@ 7.4. Handling Large AI Controllers
As mentioned in Section 5.1, the AI controllers used in our evaluations vary significantly in size, with neuron counts ranging from 440 to 1200, which are relatively large compared to those used in other studies like (Zhang et al., 2023a; Tran et al., 2020), providing a diverse basis for evaluation. However, we anticipate that larger models will be integrated into control systems as hardware continues to advance. Therefore, we conduct additional experiments to evaluate Synthify’s performance on control systems with larger AI controllers. These new experiments use three variant models trained in the CartPole, Pendulum, and Self-Driving control systems, also released by Zhu et al. (Zhu et al., 2019), which feature larger AI controllers with 3 layers containing 1200, 900, and 800 neurons, respectively—an increase of 2.4$\times$ to 6.6$\times$ the number of neurons compared to the 8 controllers used in our main experiments.
@@ 7.7. Potential Applications in End-to-End AI Systems
First and foremost, Synthify enhances the efficiency and sub-specification coverage of falsification for end-to-end AI systems. As mentioned in Section 1, these systems typically involve complex AI controllers with numerous parameters for processing inputs and generating control signals in one go (Bojarski et al., 2016; Julian et al., 2016). This complexity leads to high computational costs during testing (Haq et al., 2022; Pang et al., 2022). Synthify mitigates this by synthesizing efficient proxy programs that replace the AI controllers, significantly reducing the runtime of falsification. Moreover, thorough testing of end-to-end AI systems requires diverse test inputs (Lou et al., 2022; He et al., 2024). Synthify meets this need by maximizing sub-specification coverage of the conjunctive safety specification, as a more diverse set of test inputs is likely to falsify more sub-specifications (Zhang et al., 2023b; Viswanadha et al., 2021); intuitively vice versa. The $\epsilon$-greedy sampling strategy employed by Synthify promotes the exploration of various sub-specifications, enabling the falsification process to uncover diverse safety violations by targeting different aspects of the safety requirements.
@@
Furthermore, certain modules in Synthify, such as our synthesis and $\epsilon$-greedy sampling strategy, can complement existing verification and testing tools for end-to-end AI systems. For instance, our synthesis module could be applied to generate simple programs that closely approximate the behavior of complex AI controllers but are more amenable to formal verification methods (Fulton et al., 2020; Landers and Doryab, 2023), like constructing barrier certificates (Zhu et al., 2019) or performing reachability analysis (Tran et al., 2019). In the context of testing, our $\epsilon$-greedy sampling strategy can be integrated into existing falsification tools like (Dreossi et al., 2019a; Zhang et al., 2023a) to strengthen their ability to falsify conjunctive specifications. Alternatively, it can support automated testing frameworks that handle multiple objectives (Abdessalem et al., 2018; Haq et al., 2022), helping developers efficiently prioritize and focus on the most promising objectives during testing.
@@
Beyond testing and verification, Synthify aids developers in inspecting end-to-end AI systems for interpretability and explainability—two essential factors in ensuring the trustworthiness of AI-enabled systems (Glanois et al., 2024). Interpretability focuses on understanding the system’s inherent characteristics. Developers can leverage Synthify’s identified safety violations and corresponding falsifying inputs to derive formal specifications and properties of AI controllers (Gopinath et al., 2020; Rozier, 2016), serving as proof of the system’s safety (or lack thereof). Explainability, on the other hand, involves using external proxies to clarify the system’s operations (Vouros, 2022). Synthify’s synthesized proxy programs, with straightforward control logic, can provide explanations for generated control signals by transparently revealing the execution process. This allows developers to better understand the AI controller’s decision-making and identify potential vulnerabilities or biases within the system.
@@ 8.1. Falsification of AI-Enabled Control Systems
Falsification (Eddeland et al., 2020; Annpureddy et al., 2011; Adimoolam et al., 2017; Menghi et al., 2020; Waga, 2020; Ernst et al., 2021b; Tuncali et al., 2018; Chandratre et al., 2023) has emerged as an alternative to formal verification that provides better efficiency to uncover violations of a given safety specification. There comes many falsification techniques supported by optimization algorithms, such as Hill Climbing (Mathesen et al., 2021), Simulated Annealing (Abbas and Fainekos, 2012; Thibeault et al., 2021), Monte Carlo Tree Search (Zhang et al., 2018, 2021a), gradient-based methods (Abbas et al., 2014; Yaghoubi and Fainekos, 2017), as well as some machine learning-guided methods (Aineto et al., 2023; Zhang et al., 2022).
To date, there has been limited attention directed toward the falsification of AI-enabled control systems, with only a few early attempts (Yaghoubi and Fainekos, 2019; Song et al., 2022; Zhang et al., 2023a; Dreossi et al., 2019b; Xie et al., 2023). Among these studies, Song et al. (Song et al., 2022) evaluate the reliability of AI-enabled control systems using two existing falsification tools (Annpureddy et al., 2011; Donzé, 2010) and highlight that they need further improvements, which necessitates our work. Zhang et al. (Zhang et al., 2023a) and Yaghoubi et al. (Yaghoubi and Fainekos, 2019) develop gray-box falsification techniques that make different assumptions compared to our study. They assume the AI controllers or plants accessible and collect not only the input and output signals from the control system but also intermediate information of the AI controllers or plants as guidance for falsification. In comparison, our approach is a black-box falsification tool that solely relies on the input and output signals, making it more practical particularly in real-world applications where the controller or plants can be complex or even inaccessible if they are confidential products from third-party vendors.
@@ 8.2. Surrogate-Assisted Falsification
Another set of studies integrates surrogate models into the falsification process. For instance, several studies (Deshmukh et al., 2017; Ramezani et al., 2023; Moss et al., 2023; Zhang and Arcaini, 2021; Qin et al., 2022) employ Bayesian optimization with Gaussian processes as surrogate models to estimate the robustness semantics of STL specifications, guiding the search for safety violations. Additionally, Peltomäki et al. (Peltomäki and Porres, 2022) propose training a surrogate DNN model to directly generate falsifying inputs for control systems.
While our synthesized proxy programs may resemble surrogate models, they fundamentally differ from those in above studies. First, in terms of form, our proxy programs implement linear controllers, whereas surrogate models in prior work are typically DNNs (e.g., (Kundu et al., 2024; Peltomäki and Porres, 2022)) or machine learning models (e.g., (Deshmukh et al., 2017; Qin et al., 2022)). The simplicity of our programs offers greater efficiency. Second, regarding purpose, our proxy programs replicate the functionality of AI controllers, acting as their stand-ins to facilitate efficient falsification by minimizing the need to run the controllers themselves. This addresses the unique computational challenges associated with AI controllers, which are expensive to execute due to their vast number of parameters. In contrast, surrogate models in earlier work are usually designed to replace control systems or predict robustness semantics. Third, in terms of underlying technique, our approach utilizes Evolution Strategy for program synthesis, while prior methods often rely on Bayesian optimization (e.g., (Qin et al., 2022; Ramezani et al., 2023)) or DNNs (e.g., (Menghi et al., 2020; Kundu et al., 2024; Peltomäki and Porres, 2022)) to build surrogate models.
@@ 8.3. Program Synthesis for AI models
Although program synthesis has been studied for decades (Sobania et al., 2023), very little attention has been paid to program synthesis specialized for AI models. A few studies (Verma et al., 2018; Trivedi et al., 2024) have proposed methods for generating programmatic policies from well-trained DRL models, which are more interpretable than neural networks. Additionally, Zhu et al. (Zhu et al., 2019) introduce an inductive synthesis-based method that can synthesize deterministic and verified programs from DRL-based models. Inala et al. (Inala et al., 2020) and Qiu et al. (Qiu and Zhu, 2022) synthesize programmatic controllers to perform continuous control tasks. Our approach differs from these studies in that we propose an Evolution Strategy-based algorithm that can effectively synthesize programs as proxies for AI controllers, which we believe is both technically novel and offers a unique perspective on improving the efficiency of falsification in AI-enabled control systems.
@@ 8.4. Multi-Objective Optimization in AI-Enabled Control Systems
In the context of testing, Zolfagharian et al. (Zolfagharian et al., 2023) and Haq et al. (Haq et al., 2022, 2023) propose multi-objective search-based testing methods for AI-enabled control systems, accounting for diverse fitness functions or safety requirements. Further studies address repairing these systems. Lyu et al. (Lyu et al., 2023) repair AI controllers’ unsafe behaviors by adding patches to control signals via multi-objective optimization, while MORTAR (Wang et al., 2024b) replaces unsafe signals from AI controllers with repaired alternatives through gradient-based optimization. Both methods solve the multi-objective problem of maximizing satisfaction of safety requirements while minimizing changes to the original control signals. Regarding safety enhancement without directly repairing the violations, SIEGE (Song et al., 2023) stands out by presenting an ensemble framework that synergistically combines multiple AI controllers to improve the system’s reliability across multiple control objectives. GenSafe (Zhou et al., 2024b) corrects the behaviors of AI controllers that potentially deviate from safety constraints during training to enhance their safety while also maintaining performance at a satisfactory level. In addition to all of the above, Zhou et al. (Zhou et al., 2024a) establish a benchmark for multiple robotic manipulation tasks, providing a platform to evaluate the safety and effectiveness of AI controllers in a multi-objective context.
@@ 9. Conclusion and Future Work
Replication Package. For reproducibility, we have made our code and data publicly available: https://github.com/soarsmu/Synthify.